# PanoNormal: Monocular Indoor 360° Surface Normal Estimation

### v1

@@ 4 Experiments and Results
We adopted five popular panorama benchmark datasets for our experimental validation, both quantitatively and qualitatively. They are 3D60 [30], Stanford2D3D [3], Matterport3D [6], SunCG [20], and Structured3D [28]. In addition, SUN360 [23], which contains real-world data with no ground truth, was used for further qualitative and generalization comparison. We compared our method with HyperSphere [14] which is the state-of-the-art for 360° surface normal estimation. In addition, we further adapt four popular 360° and perspective image depth estimation methods (PanoFormer [18], OmniFusion [15], MonoViT [27], UniFuse [13]) to the surface normal estimation task, to validate the effectiveness of our designed architecture. It is essential to note that the methods compared in this evaluation represent the current state-of-the-art in both 360° surface normal estimation and a highly relevant task, depth estimation. While HRDFuse [1] is acknowledged as the latest work in 360° depth estimation which employs a histogram bin approach to learn depth distribution, it is not suitable for adapting to surface normal tasks.
@@ 4.1 Evaluation Metric and Datasets
We conducted performance evaluation in surface normal estimation using three standard angular error metrics (mean error (Mean), median error (Median), and mean square error (MSE)), and five accuracy metrics that measure the percentage of pixels where the ratio ($\delta$) between predicted surface normal vectors and ground truth is less than 5°, 7.5°, 11.5°, 22.5°, and 30°. A consistent setting is applied to ensure a fair comparison across all methods, and detailed specifics for each dataset are presented below.
4.1.1 3D60
The 3D60 dataset offers a broad spectrum of panoramic images with resolutions of $256\times 512$ captured in varied environments. The captured 360° RGB imagery with corresponding information, such as surface normal and depth with specific camera positions are from two real-world indoor scanning environments, Stanford2D3D and Matterport3D, alongside synthetic scenes from the SunCG datasets. The inherent distribution gap among these datasets enhances the model’s generalizability. To facilitate model training and evaluation, we adopt the data splits utilized in HyperSphere, as recommended in the official introduction. Notably, Matterport3D lacks ground truth data for 360° surface normals, and Stanford2D3D’s surface normal instances lack consistently aligned axes across their data. Consequently, we evaluated them based on specific separations within the 3D60 dataset.
4.1.2 Structured3D
Structured3D is a large-scale synthetic dataset featuring 21,835 $512\times 1024$ resolution of the panoramic data across 3500 scenes, encompassing RGB images illuminated with cold, normal, and warm lighting, as well as surface normal, depth, and semantic annotations. As [28] did not include any official training, validation and test splits, we preprocessed and formed the dataset examples with an 8:1:1 ratio, yielding 17,442 training data instances with three distinct lighting conditions (52,326 in total with three lighting conditions) and 2,179 and 2,181 validation and test data instances with randomly selected lighting conditions. For future convenience, the split files and the corresponding data loader are readily available in our project’s Git repository.
@@ 4.2 Implementation Details
Our experiments were carried out utilizing a single CPU core of an Intel Xeon W-2133 paired with an RTX 3090 GPU, with a batch size configured to 2 and the input resolution set to $256\times 512$. The chosen optimizer was Adam, with default settings maintained. The initial learning rate was $1\times 10^{-4}$, decreasing by half every 15 epochs. We trained our model for 110 epochs and incorporated early stopping at the 15th epoch if there were no further improvements.
@@ 4.3 Experimental Results
Table 1: Quantitative comparisons on five benchmarks with PanoNormal, HyperSphere [14], PanoFormer [18], OmniFusion [15], MonoViT [27], and UniFuse [13]. We evaluate our method improvement against the existing methods that have consistently demonstrated strong performance, distinguished by their highlighting in grey.
*We assess these datasets using the corresponding partitions of the 3D60 dataset.
@@
Ours-Improve 3.88% 25.67% 2.69% 0.79 0.80 0.55 0.28 0.19
Ours Average Improvement 9.31% 12.64% 14.01% 1.21 1.45 1.37 0.95 0.81
@@
We present a quantitative comparison among state-of-the-art methods, algorithms adapted for spherical surface normal estimation, and our PanoNormal model across five datasets, as detailed in Table 1. To ensure a fair evaluation, we retrained all models with identical settings. PanoNormal demonstrates superior performance, achieving a new state-of-the-art status across all five benchmarks. It showcases a consistent improvement of 9.31% for the mean error, 12.64% for the median error, and 14.01% for the mean square error (MSE) on average, outperforming the previous state-of-the-art methods highlighted in grey.
Figure 3: Qualitative comparisons across five benchmarks, featuring PanoNormal, UniFuse, PanoFormer, OmniFusion, MonoViT, and HyperSphere. Optimal viewing experience in color.
@@
In specific instances, our model demonstrates notable improvements in the MSE metric, achieving a 25.33% enhancement on 3D60, 19.73% on Stanford2D3D, 26.05% on Matterport3D, and 30.99% on SunCG. However, the improvement is only 2.69% on the Structured3D dataset with MSE. This difference is less significant compared to others due to the complex, synthetic nature of Structured3D scenes, which include numerous small objects with subtle curvature changes like ornaments, drawer handles, and wall-mounted kitchen tools. The dataset also presents challenges with intricate textures, as observed in items such as mirrors, glass walls, and carpets. In addition, PanoNormal exhibits a marginally higher median error compared to HyperSphere on the Stanford2D3D dataset, indicating potential sensitivity to outliers. However, this discrepancy also suggests that our method captures finer details, as HyperSphere sacrifices sharp boundaries in its predictions. The perceptibly more blurred outcomes from HyperSphere are illustrated in Fig. 3. The enhancement in $\delta$ performance underscores our model’s capacity to achieve more accurate predictions, thereby highlighting its generalizability and effectiveness across diverse domains within the datasets, including both real-world and synthetic scenarios.
@@
We also present qualitative results with the RGB input images and the corresponding predictions generated by various methods, followed by the ground truth surface normal map, illustrated in Fig. 3. We show one test example of each dataset with specific colors mapped to coordinate information, providing an intuitive representation of surface normals. Additionally, areas with unavailable data are denoted by grey on the surface normal map. The PanoNormal model stands out for its ability to capture finer details with sharper boundaries showcasing more sensitivity to the subtle surface curvatures.
@@
To assess the generalizability of our method, we conduct surface normal estimation on real-world data and then compare our results with those obtained using Hypersphere. As demonstrated in Fig. 4, PanoNormal produces precise predictions on the unseen real-world SUN360 dataset. It accurately estimates the complete boundaries of various objects, even in the presence of distortions and complex textures. Noteworthy examples include the bed at the bottom of the ERP image in the first column and the glass door in the third column. In contrast, the model of HyperSphere does not generalize well to real-world panoramic images. These observations provide further validation of the efficacy of our proposed methods. Furthermore, when examining Stanford2D3D and Matterport3D as real-world datasets, it is important to acknowledge that their ground truth surface normal map is constructed with triangle faces, and may present certain visual characteristics. However, our method demonstrates a capability for generating smooth surface normals, providing enhanced precision in defining boundaries. This improvement contributes to an elevated standard of visual quality in our results.
@@
Figure 4: Normal estimation predictions on some real-world data. The images are from the SUN360 [23] dataset. More qualitative results can be found in our supplementary materials.
@@ 4.4 Ablation Study
Table 2: Ablation study.
Method PanoFormer PF-R1 Ours Ours-R1 Ours-R2 Ours-R3 Ours-R4
Mean $\downarrow$ 17.500 5.154 4.931 5.310 5.337 5.305 5.506
Median $\downarrow$ 6.584 0.321 0.267 0.394 0.383 0.358 0.398
MSE $\downarrow$ 929.233 178.146 167.717 183.356 186.472 184.613 193.994
In our comprehensive evaluation under identical conditions, we conducted an ablation study on the 3D60 dataset to validate the key components of our model. As presented in Table 2, PanoFormer serves as the baseline structure for surface normal prediction, adapted from depth estimation methods. Introducing the multi-level decoder in the first placement of PanoFormer results in a notable decrease in mean, median, and MSE error metrics by 70.55%, 95.12%, and 80.83%, respectively. Our network, featuring another proposed feature embedding component, emerges as the top-performing configuration among ablation experiments, achieving state-of-the-art performance by another improvement with 4.32% on the mean, 16.84% on the median, and 5.85% on MSE. The "Replace 1 – 4" scenario, where convolutional layers replace self-attention blocks symmetrically in both the encoder and decoder, demonstrates promising results. Specifically, the third placement, maintaining the self-attention block only at the bottleneck and one level higher in both the encoder and decoder, achieves the second-best performance on average. This configuration holds potential for real-world applications with limited device resources, offering advantages in terms of reduced memory space, faster training, and quicker inferencing times. Fig. 5 illustrates qualitative comparisons among our ablation architectures. In the first row, we display the input panoramic images along with their corresponding ground truth surface normals. The following rows depict the normalized absolute error between the predicted surface normal map and the ground truth map. In these rows, darker shades of red indicate higher error, alongside the corresponding predicted surface normal maps. Once more, it is evident that our current architecture exhibits superior performance when confronted with diverse challenging scenarios.
Furthermore, our proposed architecture consistently outperforms PanoFormer  [18] across all metrics, both quantitatively and qualitatively, on every dataset. The adapted PanoFormer implementation yields less satisfactory results in all comparisons, encountering notable challenges on the Structured3D synthetic dataset and struggling with convergence. In contrast, our PanoNormal framework consistently demonstrates outstanding performance across all datasets, highlighting the efficacy and significance of the proposed architecture.
Figure 5: Qualitative comparisons among our ablation architectures. Top row: RGB inputs and the ground truth surface normal maps. Subsequent rows: the experiment architectures with predicted surface normals and corresponding error maps, where darker shades of red signify more pronounced errors. Optimal viewing experience in color.
4.5 Limitation
@@
In Table 1, our method exhibits a performance of approximately 24.59% higher median error compared to HyperSphere on the Stanford2D3D dataset while demonstrating significant improvements in the metrics of mean and MSE. The enhancements on mean and MSE suggest that, on average, our model provides more accurate predictions. However, the higher median error implies the existence of a subset of predictions with larger errors.
@@
Table 3: Specific examples from Stanford2D3D dataset.
Example Method Mean $\downarrow$ Median $\downarrow$ MSE $\downarrow$ $\delta<$ 7.5° $\delta<$ 22.5° $\delta<$ 30°
(a) HyperSphere 10.5388 0.0656 470.1378 66.24 82.44 86.18
Ours 10.4650 0.0816 450.3420 66.57 82.43 86.70
(b) HyperSphere 16.8774 7.6914 753.1895 49.69 68.10 75.56
Ours 16.1840 8.8487 634.5417 47.76 68.10 76.95
(c) HyperSphere 13.2489 0.8248 564.9490 57.08 77.85 83.39
Ours 12.2771 2.3333 473.4669 58.34 79.36 84.92
Figure 6: Qualitative evaluation on three specific examples.

### v2

@@ 4 Experiments and Results
We conducted experimental validation using five widely recognized panorama datasets: 3D60 [zioulis2018omnidepth], Stanford2D3D [armeni2017joint], Matterport3D [chang2017matterport3d], SunCG [song2017semantic], and Structured3D [zheng2020structured3d]. These datasets were used for both quantitative and qualitative analysis. Additionally, we included the SUN360 [xiao2012recognizing] dataset, which contains real-world data without ground truth, for further qualitative and generalization comparisons.
To evaluate the performance of our method, we compared it against HyperSphere [karakottas2019360], which is the current state-of-the-art method specifically designed for 360° surface normal estimation, and 360MTL [huang2025multi], which predicts 360° surface normals as a sub-task. We also adapted five popular 360° and perspective image depth estimation methods—PanoFormer [shen2022panoformer], OmniFusion [li2022omnifusion], MonoViT [zhao2022monovit], and UniFuse [jiang2021unifuse]—by modifying their final prediction layer to output three dimensions instead of one, thereby making them applicable to the surface normal estimation task. This adaptation was crucial for validating the effectiveness of our proposed architecture.
@@ 4.1 Evaluation Metric and Datasets
We conducted performance evaluation of surface normal estimation using three standard angular error metrics (mean error (Mean), median error (Median), and mean square error (MSE)), and five accuracy metrics that measure the percentage of pixels where the ratio ($\delta$) between the predicted surface normal vector and the ground truth is less than 5°, 7.5°, 11.5°, 22.5°, and 30°. A consistent setting is applied to ensure a fair comparison across all methods, and detailed specifics for each dataset are presented as follows.
3D60 dataset offers a broad spectrum of panoramic images with resolutions of $256\times 512$ captured in varied environments. The captured 360° RGB imagery with corresponding information, such as surface normal and depth with specific camera positions are from two real-world indoor scanning environments, Stanford2D3D and Matterport3D, alongside synthetic scenes from the SunCG datasets. The inherent distribution gap among these datasets enhances the model’s generalizability. To facilitate model training and evaluation, we adopt the data splits utilized in HyperSphere, as recommended in the dataset’s documentation. Notably, Matterport3D lacks ground truth data for 360° surface normals, and Stanford2D3D’s surface normal instances lack consistently aligned axes across their data. Consequently, we evaluated them based on specific separations within the 3D60 dataset.
Structured3D is a large-scale synthetic dataset featuring 21,835 RGB images of resolution $512\times 1024$. These panoramic images were captured across 3500 scenes, illuminated with cold, normal, and warm lighting. The dataset also includes surface normal, depth, and semantic annotations. We preprocessed and formed the dataset examples with an 8:1:1 ratio, yielding 17,442 training data instances with three distinct lighting conditions (52,326 in total with three lighting conditions) and 2,179 and 2,181 validation and test data instances with randomly selected lighting conditions.
@@ 4.2 Implementation Details
Our experiments were carried out using a single CPU core of an Intel Xeon W-2133 paired with an RTX 3090 GPU, with a batch size configured to 2 and the input resolution set to $256\times 512$. The chosen optimizer was Adam, with default settings maintained. The initial learning rate was $1\times 10^{-4}$, decreasing by half every 15 epochs. We trained our model for 110 epochs and incorporated early stopping at the 15th epoch if there were no further improvements.
@@ 4.3 Experimental Results
Table 1: Quantitative comparisons. *Evaluated on corresponding splits of the 3D60 dataset.
@@
Ours-Improve 3.88% –4.28% 2.69% 0.79 0.80 0.55 0.28 0.19
Ours Average Improvement 6.50% –2.34% 8.96% 0.81 0.84 0.75 0.59 0.51
@@
We present a comprehensive quantitative comparison between state-of-the-art methods (HyperSphere and 360MTL), algorithms adapted for spherical surface normal estimation, and our PanoNormal model across five datasets, as detailed in Table 1. For a fair evaluation, all models were retrained under identical settings. PanoNormal demonstrates superior performance, achieving state-of-the-art results across all five benchmarks. On average, it delivers a consistent improvement of 6.50% in mean error and 8.96% in MSE, surpassing the previous best methods highlighted in grey.
In specific cases, our model shows significant improvements in the MSE metric: 15.22% on 3D60, 10.94% on Stanford2D3D, 14.81% on Matterport3D, and 25.10% on SunCG. However, the improvement is relatively modest on the Structured3D dataset, with only a 2.69% enhancement in MSE. This less significant difference is due to the complex, synthetic nature of Structured3D scenes, which contain numerous small objects with subtle curvature changes, such as ornaments, drawer handles, and wall-mounted kitchen tools. Additionally, the dataset challenges the models with detailed textures in items like mirrors, glass walls, and carpets.
@@
In addition, PanoNormal shows a marginally higher median error than HyperSphere on the Stanford2D3D dataset, indicating potential sensitivity to outliers. However, this discrepancy also suggests that our method captures finer details. HyperSphere, on the other hand, sacrifices sharp boundaries in its predictions, resulting in more blurred outcomes, as illustrated in Fig. 3.The improvement in $\delta$ performance underscores PanoNormal’s ability to achieve more accurate predictions, demonstrating its generalizability and effectiveness across diverse domains, including both real-world and synthetic scenarios within the datasets.
@@
We also provide qualitative results, showcasing RGB input images alongside the corresponding predictions generated by various methods, with the ground truth surface normal map shown in Fig. 3. Each dataset’s test example is displayed with specific colors mapped to coordinate information, offering an intuitive representation of surface normals. Areas with unavailable data are denoted by grey on the surface normal map. The PanoNormal model stands out for capturing finer details and sharper boundaries, showing greater sensitivity to subtle surface curvatures.
Figure 3: Qualitative comparisons across five benchmarks, featuring PanoNormal, UniFuse, PanoFormer, OmniFusion, MonoViT, HyperSphere, and 360MTL. Optimal viewing experience in color.
@@
To assess the generalizability of our method, we conduct surface normal estimation on real-world data, comparing our results with those from HyperSphere and 360MTL. As demonstrated in Fig. 4, PanoNormal produces precise predictions on the unseen real-world SUN360 dataset, accurately estimating the complete boundaries of various objects even in the presence of distortions and complex textures. Noteworthy examples include the bed at the bottom of the ERP image in the first row and the glass door in the third row. In contrast, HyperSphere does not generalize well to real-world panoramic images, further validating the efficacy of our proposed method.
When examining Stanford2D3D and Matterport3D as real-world datasets, it is important to note that their ground truth surface normal maps are constructed with triangle faces, which may introduce certain visual characteristics. However, our method shows a strong ability to generate smooth surface normals with enhanced boundary precision, contributing to a higher standard of visual quality in our results. More discussion are available in the supplementary material.
@@
Figure 4: Normal estimation predictions on some real-world data. The images are from the SUN360 [xiao2012recognizing] dataset. More qualitative results can be found in our supplementary materials.
To evaluate the effectiveness of applying perspective image algorithms to the 360° domain, we conduct experiments using ASNGeo [long2024adaptive], a recent multi-task model designed to predict both depth and surface normal maps for perspective images. ASNGeo includes a depth-to-normal (D2N) layer that calculates surface normals by averaging directions from multiple randomly sampled triangle faces covering each 3D point. We retrained ASNGeo using 360° data, keeping the default settings for the D2N layer to maintain consistency with the original paper. As shown in Fig. 5, the qualitative results from ASNGeo’s surface normal prediction branch reveal suboptimal performance, even when objects in the ERP image exhibit minimal distortion. We also apply the D2N layer to depth maps from both ASNGeo and a recent 360° depth estimation approach (GLPanoDepth [bai2024glpanodepth]). The results show blurry outputs around objects and inconsistent appearances for ceilings and floors. Additionally, deriving surface normals from depth using this method accumulates errors, ultimately leading to poorer results. These highlight the necessity of developing specialized algorithms for 360° surface normal estimation.
Figure 5: Surface normal estimation predictions on Stanford2D3D data are compared between ASNGeo [long2024adaptive] (a perspective-based method) and our approach. To further validate the conversion process, we used the depth-to-surface-normal technique from ASNGeo [long2024adaptive] to convert both perspective depth maps and GLPanoDepth [bai2024glpanodepth] (360° domain) depth maps into surface normals.
@@ 4.4 Ablation Study
Table 2: Ablation study for individual components. Beginning with the PanoFormer baseline, which has no feature embedding and a multi-level decoder, we incrementally add each component one at a time.
Method Mean $\downarrow$ Median $\downarrow$ MSE $\downarrow$ Baseline (B) 17.500 6.584 929.233 B + multi-level decoder 5.154 0.321 178.146 B + FeatEmbed + multi-level decoder 4.931 0.267 167.717
4.4.1 Individual component study.
We conducted an ablation study to validate the key components of our model. As presented in Tab. 2, PanoFormer serves as the baseline structure for surface normal prediction. Originally designed for depth estimation, PanoFormer exhibits suboptimal performance when directly applied to surface normals, primarily due to the distinct nature of the tasks. Surface normal estimation relies on fine-grained local geometric details, whereas depth estimation emphasizes global scene structure, highlighting the architectural mismatch when the baseline is used without modification.
To address these limitations we introduce, first, a multi-level decoder, and then add a feature embedding module. The multi-level decoder was particularly impactful, resulting in significant reductions in mean, median, and MSE error metrics by 70.55%, 95.12%, and 80.83%, respectively. We believe that its effectiveness arises from its ability to aggregate multi-scale information, combining low-level features for detailed geometry with high-level contextual cues, which is crucial for accurately predicting surface normals across diverse regions, including areas with sharp boundaries or fine textures. This capability ensures that both global consistency and local precision are maintained, addressing one of the primary shortcomings of the baseline.
Building on this, the feature embedding module was designed to enhance the representation of fine local geometric structures, capturing intricate surface characteristics such as subtle curvature changes and texture discontinuities that are often overlooked in global feature representations. Integrating this module into the architecture led to further improvements, with reductions of 4.32% in the mean error, 16.84% in the median error, and 5.85% in MSE. The module effectively complements the multi-level decoder by focusing on local feature refinement, enabling the model to perceive nuanced geometric variations in complex scenes.
Together, these components address three critical factors: the task-specific mismatch in the baseline architecture, the need for effective multi-scale feature aggregation, and the importance of capturing detailed local geometry. By leveraging these enhancements, our model achieves a significant improvement in surface normal prediction accuracy, demonstrating the value of balancing global context with fine-grained detail. These findings underscore the necessity of architectural adaptations for domain-specific tasks like surface normal estimation and the potential of our approach to set a new benchmark in this field.
Table 3: Ablation study for loss combinations. Starting with the MSE and perceptual loss, we progressively incorporate the quaternion and smooth loss.
Loss Mean $\downarrow$ Median $\downarrow$ MSE $\downarrow$ $L_{m}$ + $L_{p}$ 5.277 0.593 170.768 $L_{m}$ + $L_{p}$ + $L_{q}$ 5.130 0.314 179.481 $L_{m}$ + $L_{p}$ + $L_{s}$ 5.421 0.660 174.738 $L_{m}$ + $L_{p}$ + $L_{q}$ + $L_{s}$ 4.931 0.267 167.717
Table 4: Ablation study evaluating the effect of varying the number of ViT and convolutional blocks in our model. We also report comparison with HyperSphere.
Method #\#ViT block #\#Conv block GFLOPs $\downarrow$ Params (M) $\downarrow$ FPS $\uparrow$ Mean $\downarrow$ Median $\downarrow$ MSE $\downarrow$ Ours 9 0 179.60 35.14 10.13 4.931 0.267 167.717 Ours-V1 7 2 172.04 35.09 20.65 5.310 0.394 183.356 Ours-V2 5 4 185.50 34.98 32.19 5.337 0.383 186.472 Ours-V3 3 6 211.80 34.72 41.30 5.305 0.358 184.613 Ours-V4 1 8 223.68 30.16 53.39 5.506 0.398 193.994 Ours-V5 0 9 224.50 30.98 60.81 5.533 0.360 202.060 HyperSphere - - 197.94 35.12 68.41 5.784 0.266 224.623
4.4.2 Loss combinations.
We investigate the effectiveness of various losses on our model’s accuracy to determine the optimal combination for training. As shown in Tab. 3, we began with the common losses: MSE ($L_{m}$) and perceptual loss ($L_{p}$), on the 3D60 dataset. Adding the quaternion loss ($L_{q}$) improved the mean and median error metrics by 2.79% and 47.05%, respectively, though it increased the MSE error, indicating higher sensitivity to outliers. Conversely, adding the smooth loss ($L_{s}$) alone decreased performance. However, combining all four losses with specific weights yielded the best results. This suggests that while the quaternion loss sharpens scene details but is sensitive to outliers, incorporating smooth loss helps achieve a balance, enhancing fine-grained details without sensitivity to outliers.
4.4.3 Number of ViT blocks
The number of ViT blocks affects both the accuracy and efficiency of our method. To investigate this, we progressively replaced pairs of self-attention blocks with convolutional blocks in both the encoder and decoder, resulting in multiple configurations with different ViT-to-convolution ratios (Tab. 4). The results show that accuracy consistently decreases as more convolutional layers are substituted, confirming the effectiveness of ViT blocks for capturing long-range dependencies. The best performance is obtained when all blocks are ViTs. Interestingly, the variant “Ours-V3,” which retains self-attention only at the bottleneck and one higher level in both encoder and decoder, achieves competitive accuracy while significantly simplifying the architecture. This indicates that a small number of strategically placed ViT blocks can still preserve most of the benefits of global reasoning, while potentially making the model more suitable for deployment on resource-constrained devices.
4.4.4 Efficiency of our models
In addition to accuracy, we report computational cost (GFLOPs), model size, and inference speed (frame per second, FPS) in Tab. 4. These results highlight a clear trade-off between accuracy and efficiency: ViT-heavy configurations deliver the best error metrics, whereas conv-heavy variants run faster and require fewer resources. This flexibility allows our framework to be adapted to different application requirements, such as prioritizing accuracy in offline high-fidelity settings or emphasizing speed in real-time scenarios. Compared to HyperSphere, our models offer a more favorable balance, achieving comparable or better accuracy with competitive computational efficiency. For instance, mid-range variants like Ours-V3 provide a good compromise, maintaining strong predictive performance while reducing memory usage and improving inference speed, making them attractive candidates for practical deployment.
5 Limitation
@@
In Tab. 1, our method exhibits a performance of approximately 24.59% higher median error compared to HyperSphere on the Stanford2D3D dataset while demonstrating significant improvements in the metrics of mean and MSE. The enhancements on mean and MSE suggest that, on average, our model provides more accurate predictions. However, the higher median error implies the existence of a subset of predictions with larger errors.
Table 5: Specific examples from Stanford2D3D dataset.
Example Method Mean $\downarrow$ Median $\downarrow$ MSE $\downarrow$ (a) HyperSphere 10.5388 0.0656 470.1378 Ours 10.4650 0.0816 450.3420 (b) HyperSphere 16.8774 7.6914 753.1895 Ours 16.1840 8.8487 634.5417 (c) HyperSphere 13.2489 0.8248 564.9490 Ours 12.2771 2.3333 473.4669
Figure 6: Qualitative evaluation on three specific examples. Note that the Ground Truth is based on triangular faces, which is not realistic. HyperSphere achieves lower median error but blurs object boundaries, while our method predicts more precise boundaries that align better with the actual shape of objects.
@@
6 More Comparison on SUN360
Figure 7: Qualitative comparison of the models for the SUN360 dataset. Please note that the SUN360 dataset is a real-world dataset and does not include ground truth data. We provide qualitative comparison of the models across the entire SUN360 which can be found in the separated supplementary materials.
In Fig. 7, we show qualitative comparisons on the SUN360 dataset for 360 surface normal estimation. Our model demonstrates superior performance compared to existing state-of-the-art approaches, includes models such as HyperPSphere, PanoFormer, MonoViT, UniFuse, and OmniFusion, all of which have been evaluated on diverse indoor scenes.
Our model consistently delivers more accurate and coherent surface normal predictions, as evident from the sharpness and continuity of the estimated surfaces across varying room structures. Unlike competing models, which often produce fragmented or blurred normal maps, especially in areas with complex geometries, our approach maintains clarity and precise border of surfaces. This is particularly noticeable in regions with intricate details, such as furniture edges and room corners, where other methods tend to falter.
These results highlight the robustness and effectiveness of our method in handling the challenges posed by real-world 360° imagery. Our method excels in maintaining visual consistency and preserving fine details across various indoor environments, outperforming other models in the process. The ability of our model to produce clear and accurate surface normal estimations, even in complex scenes, sets a new standard for 360 surface normal estimation and establishes it as a reliable solution for applications that demand a high level of environmental understanding.